The sum of the binomial coefficients of is equal to 256 . The constant term in the expansion is
1120
2110
1210
none

Grade:12th Pass

1 Answers

Nehal Wani
13 years ago

[2x+1/x]n

Sum of binomial coefficients:

nC0+ nC1+ nC2+ nC3+………………….+ nCn+ nCn=2n

256=2n

Therefore n=8

Tr=nCr(2x)n-r(x)-r= nCr2n-rxn-2r= 8Cr28-rx8-2r

8-2r=0

⇒  r=4

8Cr28-rx8=8C424=1120